OpenStudy (anonymous):

Write each expression in terms of sin(theta), cos(theta), or both. Simplify wherever possible. Its trig identities 1. tan(theta) * csc(theta) = ??? 2. tan(theta) / sec(theta) = ???

OpenStudy (jdoe0001):

http://www.mathwords.com/t/trig_identities.htm

OpenStudy (anonymous):

I know what the identities are I just don't know how to solve the problem

OpenStudy (jdoe0001):

\(\bf tan(\theta)\cdot csc(\theta)\implies \cfrac{\cancel{sin(\theta)}}{cos(\theta)}\cdot \cfrac{1}{\cancel{sin(\theta)}}\)

OpenStudy (anonymous):

so its \[1 / \cos(\theta)\] ?

OpenStudy (jdoe0001):

yes, and what's \(\bf \cfrac{1}{cos(\theta)} \quad ?\)

OpenStudy (anonymous):

or sec(theta)

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

but is that still in terms of sin or cos (theta) ??

OpenStudy (jdoe0001):

hmmm ohhh yeah... forgot that part... so ... .leave it like \(\bf \cfrac{1}{cos(\theta)}\) then

OpenStudy (anonymous):

okay thanks a lot

OpenStudy (jdoe0001):

\(\bf \cfrac{tan(\theta)}{sec(\theta)}\implies \cfrac{ \frac{sin(\theta)}{cos(\theta)} }{ \frac{1}{cos(\theta)} }\implies \cfrac{sin(\theta)}{\cancel{cos(\theta)}} \cdot \cfrac{\cancel{cos(\theta)}}{1}\)

OpenStudy (jdoe0001):

yw